# Unmasking Social Bots: How Confident Are We?

### v1

@@ Results
$|P_{pred}-0.5|>3\sigma(P_{pred}).$ (1)
Equation (1) is used for both $\sigma(P_{\text{pred}})=\sigma_{\text{epi.}}(P_{\text{pred}})$ and $\sigma_{\text{tot.}}(P_{\text{pred}})$, representing the cases of epistemic only or the total uncertainty in quadrature between aleatoric and epistemic. The results show an improvement in performance, across all metrics, over the baseline as uncertainty is introduced into the decision making process. We also report the “rejection” fraction, which corresponds to the number of account that are held for further information to be acquired before classification. This additional information from uncertainty allows for more informed decisions, thereby impacting the decision-making process for Twitter/X accounts, as further discussed below.
Human Bot
BLOC Baseline 92.474 $\pm$ 0.003 0.93 0.92 0.92 0 0.92 0.93 0.93 0
Epistemic 95.836 $\pm$ 0.002 0.96 0.96 0.96 10.6 0.96 0.96 0.96 12.2
Aleatoric 96.027 $\pm$ 0.002 0.96 0.96 0.96 12.2 0.96 0.96 0.96 14.0
Quadrature 96.608 $\pm$ 0.002 0.96 0.97 0.97 15.5 0.97 0.97 0.97 17.6
Botometer Baseline 93.863 $\pm$ 0.002 0.94 0.94 0.94 0 0.94 0.94 0.94 0
Epistemic 96.537 $\pm$ 0.002 0.96 0.97 0.97 7.3 0.97 0.96 0.97 7.5
Aleatoric 96.541 $\pm$ 0.002 0.96 0.97 0.97 7.8 0.97 0.96 0.97 10.0
Quadrature 97.172 $\pm$ 0.002 0.97 0.98 0.97 10.3 0.98 0.97 0.97 12.1
Table 2: Performance Comparison with $3\sigma$ Uncertainty Thresholds: Evaluation of the uncertainty informed decision making process, in which a $3\sigma$ cut is applied on the probability around 0.5 to indicate whether decisions should be made about a user account. $3\sigma$ cuts are applied to the epistemic, aleatoric and uncertainties in quadrature. Note the increase in performance as specific accounts are withheld. Extra information from these accounts can be accquired to reduce uncertainty to a desirable threshold, or allow human intervention on more reasonable sample sizes.
Discussion
Our study details a fully Bayesian framework inspired by ELUQuant [38] to classify Twitter/X accounts as bots or humans, assessing its performance against DNNs and RF models. The BNN demonstrates comparable performance to both the DNN and RF models in terms of AUC, while providing additional information in the form of uncertainty. Crucial for decision-making at the account level. Closure tests affirm the robustness of our uncertainty quantification, in which we are able to decouple the aleatoric and epistemic components. By applying a $3\sigma$ uncertainty threshold, we observe improved accuracy and F1 scores, underscoring the utility of uncertainty-aware models in bot detection. This approach not only enhances predictive reliability but also provides deeper insights into model behavior, advancing the field’s understanding and application of probabilistic neural networks.
Our contribution should be framed in the context of end-to-end analysis pipelines using uncertainty at the account level. As we have shown, our design philosophy is agnostic to input, obtaining similar performance on both BLOC and Botometer features. The network itself can easily be adapted to more complex problems and deploy the same uncertainty aware procedures developed within. Specifically, using uncertainty to isolate subsets of accounts where the network has shown to be unreliable. These accounts can be further monitored over a period of time and reevaluated once more information has been obtained. Reducing the potential of false account suspension.
Methods
Bayesian Neural Networks
Bayesian Neural Networks (BNNs) are extensions of traditional Deep Neural Networks (DNNs), aiming to optimize distributions of weights at each layer. [^1] In contrast to their deterministic counterparts, which define a fixed set of weights over the model. The aim of BNNs is to approximate a posterior distribution over the weights, given a dataset $q(\mathbb{W}|D)$. This allows predictions of quantities through a posterior distribution $q(\mathbb{y}|\mathbb{x},D)$, integrated over the weights. The formulation of such a posterior is intractable and therefore one must turn to Bayesian inference techniques during optimization. Traditional approaches define the posterior distribution to be a fully factorized Gaussian (diagonal covariance) $q(\mathbb{W})$, such that the evidence-lower bound (ELBO) can be maximized between the approximated posterior and the prior. Note that maximizing the ELBO is equivalent to minimizing the Kullback-Leibler (KL) Divergence. This assumption, although more advantageous, tends to be limiting in terms of learned network complexity. Another approach provided in Louizos and Welling [40], is to instead approximate the posterior as a product of fully factorized Gaussian and a mixing density, Eq. 2, where $q(\mathbb{z})$ is a vector of random variables. The random variables act multiplicatively on the means of the mixing density to reduce computational complexity.
[…]
@@ Methods
The result is a more flexible posterior distribution over the weights, capable of learning multi-modal dependencies between weights. However, this too becomes intractable, and therefore, an approximate lower bound of entropy must be constructed. This can be done through an auxiliary distribution $r(\mathbb{z}|\mathbb{W})$, equivalent to performing variational inference on an augmented probability space [40]. Moreover, the approximated posterior remains true given the fact that the auxiliary distribution can be marginalized out. As stated in Louizos and Welling [40], the tightness of the bound on $q(\mathbb{W})$ (and therefore the quality of it) directly depends on the ability of $r(\mathbb{z}|\mathbb{W})$ to approximate the posterior of $q(\mathbb{z}|\mathbb{W})$, and is therefore chosen to be represented with inverse normalizing flows. The choice of normalizing flows allows analytic computation of the marginals through bijective transformations. The approximate posterior is then constrained through Eq. Bayesian Neural Networks during training, acting as a regularization term in conjunction with traditional loss functions.
@@
We also extend our network to capture the aleatoric uncertainty component using the methodology described in Kendal et al. [41]. For each input the network produces a latent variable $\mathbb{f}$, along with with the aleatoric uncertainty component $\mathbb{s}=\log\sigma$. We choose to interpret the network output ($\mathbb{s}$) as the logarithm of the uncertainty, allowing $\sigma$ to be positive definite through an exponential transformation $\sigma=e^{\mathbb{s}}$. We then define a Gaussian distribution over the latent variate such that:
@@
As mentioned in Kendal et al. [41], it is not possible to integrate out the Gaussian distribution, and as such Monte Carlo integration must be deployed. At training time this amounts to and extra sampling step to draw samples following Eq. 6 in which we take the expected value of the latent variable to produce our probability.
@@
We inherit the design philosophy from Fanelli and Giroux [38], in which we first design a minimal complexity DNN as the basis for our BNN. Bayesian blocks characterized by Multiplicative Normalizing Flows (MNF) layers [40] are utilized at each layer. The analysis pipeline is shown in Fig. 4.
Figure 4: Analysis Pipeline: Schematic representation of uncertainty aware decision making in bot detection. The Bayesian Neural Network (BNN) structure is characterized by Multiplicative Normalizing Flows (MNF) [40], batch normalization, and SELU activation functions [42]. The output of the network is the probability of a bot account, along with the epistemic and aleatoric uncertainties. These uncertainties can be combined in quadrature.
As stated in Fanelli and Giroux [38], SELU activation functions, as presented by Klambauer et al. [42], possess inherent self-normalizing properties, which ensure non-vanishing gradients. Their self-normalization nature could provide cases in which batch normalization is not needed, although this is data-dependent. We utilize SELU along with batch normalization to improve network convergence [42]. The output of the network provides a probability of the account being a bot given a BLOC input; a representation of both the accounts actions and content. The task is treated as a binary classification loss where we deploy Binary Cross-Entropy, Eq. 7, coupled to the KL term in Eq. Bayesian Neural Networks.
@@
The network is trained using a traditional $70/15/15\%$ split, in which we make the distribution of humans and bots equal prior to splitting. This removes any bias the network may incur due to class imbalance. This sampling results in an excess of human accounts which we use as additional performance measures. The Adam optimizer is used, along with a Cosine Annealing learning rate scheduler with an initial learning rate of $5\times 10^{-4}$. We deploy early stopping, defining the convergence when the validation loss is no longer decreasing after five epochs. The number of epochs for early stopping is chosen to reflect the stability of the training and account for fluctuations. Information regarding training is summarized in Table. 3.
@@
BLOC
The Behavioral Languages for Online Characterization (BLOC [35]) provides formal languages that represent the behaviors of social media accounts irrespective of social media platform, user-agent (human or bot), or intent (malicious or benign). The BLOC formal languages are defined by a set of alphabets (action and content) and rules for generating BLOC strings which are tokenized to produce BLOC words. BLOC words which represent the behaviors of social media accounts, consist of symbols drawn from distinct alphabets representing an account’s actions and content. Fig. 5a illustrates a possible representation of a sequence of tweets by three different Twitter accounts, @NASA, @Alice, and @Bob. The @NASA account replied to a tweet, posted a tweet, and then re-shared (retweet) a tweet, resulting in the BLOC action sequence: $p.T.r$. Here, each action (e.g., reply to) is represented by a single symbol (e.g., $p$), and the dots represent long pauses (e.g., $>$ 1 minute) between actions.
Once generated, BLOC strings are be tokenized (Fig. 5b) into bi-grams (two-letter words), then we can represent any social media account as a vector of BLOC words. A collection of point vectors corresponding to multiple accounts make up the BLOC matrix in Fig. 5c. In this vector space model, each account is represented as a point $(w_{1},w_{2},...,w_{k})$ in $k$-dimensional vector space where each dimension $i$ corresponds to a BLOC word. The weight $w_{i}$ represents how well an account is described by word $i$. For each account $a$, we instantiated $w_{i}$ with the TF-IDF weight [43], the product of the term frequency (TF) and the inverse document frequency (IDF):
$w_{i}(a)=f_{i}(a)\left(1+\log\frac{D}{d_{i}}\right)$ (9)
Botometer
Botometer,222botometer.org is a publicly available supervised machine learning system that classifies a given Twitter account as bot or human-controlled. Botometer-V4 (the current version of Botometer at the time of writing) [27] utilizes over 1,000 features that can be grouped into six categories that focus on different account characteristics including, metadata from the accounts (e.g., numbers of friends and followers), retweet and mention networks, temporal features (e.g., frequency of posts), content information, and sentiment. In the deployed system, different classifiers in an ensemble are trained on different accounts types, and then these classifiers vote to obtain the final bot score [27]. Here however, we only utilize the representation power of Botometer by extracting its features which serve as input to BNN and DNN.
Conclusion

### v2

@@ 3.2 Bayesian Neural Networks
The result is a more flexible posterior distribution over the weights, capable of learning multi-modal dependencies between weights. However, this too becomes intractable, and therefore, an approximate lower bound of entropy must be constructed. This can be done through an auxiliary distribution $r(\mathbb{z}|\mathbb{W})$, equivalent to performing variational inference on an augmented probability space Louizos and Welling (2017). Moreover, the approximated posterior remains true given the fact that the auxiliary distribution can be marginalized out. As stated in Louizos and Welling Louizos and Welling (2017), the tightness of the bound on $q(\mathbb{W})$ (and therefore the quality of it) directly depends on the ability of $r(\mathbb{z}|\mathbb{W})$ to approximate the posterior of $q(\mathbb{z}|\mathbb{W})$, and is therefore chosen to be represented with inverse normalizing flows. The choice of normalizing flows allows analytic computation of the marginals through bijective transformations. The approximate posterior is then constrained through Eq. 3.2 during training, acting as a regularization term in conjunction with traditional loss functions.
@@
We also extend our network to capture the aleatoric uncertainty component using the methodology described in Kendal et al. Kendall and Gal (2017). For each input the network produces a latent variable $\mathbb{f}$, along with with the aleatoric uncertainty component $\mathbb{s}=\log\sigma$. We choose to interpret the network output ($\mathbb{s}$) as the logarithm of the uncertainty, allowing $\sigma$ to be positive definite through an exponential transformation $\sigma=e^{\mathbb{s}}$. We then define a Gaussian distribution over the latent variate such that:
@@
As mentioned in Kendal et al. Kendall and Gal (2017), it is not possible to integrate out the Gaussian distribution, and as such Monte Carlo integration must be deployed. At training time this amounts to and extra sampling step to draw samples following Eq. 6 in which we take the expected value of the latent variable to produce our probability.
@@
We inherit the design philosophy from Fanelli and Giroux Fanelli and Giroux (2024), in which we first design a minimal complexity DNN as the basis for our BNN. Bayesian blocks characterized by Multiplicative Normalizing Flows (MNF) layers Louizos and Welling (2017) are utilized at each layer. The analysis pipeline is shown in Fig. 3.
Figure 3: Analysis Pipeline: Schematic representation of uncertainty aware decision making in bot detection. The Bayesian Neural Network (BNN) structure is characterized by Multiplicative Normalizing Flows (MNF) Louizos and Welling (2017), batch normalization, and SELU activation functions Klambauer et al. (2017). The output of the network is the probability of a bot account, along with the epistemic and aleatoric uncertainties. These uncertainties can be combined in quadrature.
As stated in Fanelli and Giroux Fanelli and Giroux (2024), SELU activation functions, as presented by Klambauer et al. Klambauer et al. (2017), possess inherent self-normalizing properties, which ensure non-vanishing gradients. Their self-normalization nature could provide cases in which batch normalization is not needed, although this is data-dependent. We utilize SELU along with batch normalization to improve network convergence Klambauer et al. (2017). The output of the network provides a probability of the account being a bot given a BLOC input; a representation of both the accounts actions and content. The task is treated as a binary classification loss where we deploy Binary Cross-Entropy, Eq. 7, coupled to the KL term in Eq. 3.2.
@@
4 Datasets and Experimental Setup
For this study, we utilized a set of datasets from the Bot Repository, [^3] https://botometer.osome.iu.edu/bot-repository which consists of labeled Twitter account data gathered by various researchers between 2017 and 2019. These datasets were specifically created to assist in the development of bot detection models and encompass accounts from diverse domains, including political discourse, celebrity interactions, and general social media activity. In total, the dataset includes over 10 million tweets. The datasets used in this study, along with the number of labeled bot and human accounts, are presented in Table 1.
Source Bot Counts Human Counts
midterm-18 Yang et al. (2020) 0 7458
cresci-stock-18 Cresci et al. (2018) 7102 6172
pronbots-19 Yang et al. (2019) 17884 0
botometer-feedback-19 Yang et al. (2019) 139 379
cresci-17 Cresci et al. (2017b) 7049 2760
celebrity-19 Yang et al. (2019) 0 20549
gilani-17 Gilani et al. (2017) 914 1576
verified-19 Yang et al. (2020) 0 1891
astroturf-20 Sayyadiharikandeh et al. (2020) 502 0
vendor-purchased-19 Yang et al. (2019) 1069 17
varol-17 Varol et al. (2017) 732 1496
political-bots-19 Yang et al. (2019) 62 0
cresci-rtbust-19 Mazza et al. (2019) 352 340
botwiki-19 Yang et al. (2020) 691 0
Total 36496 42638
Table 1: Bot and Human Distribution across Source Datasets: Account distribution across source datasets used within this study. The accounts are extracted from individual datasets, and then combined to form our final dataset.
The BNN is trained using a traditional $70/15/15\%$ split, in which we make the distribution of humans and bots equal prior to splitting. This removes any bias the network may incur due to class imbalance. This sampling results in an excess of human accounts which we use as additional performance measures. The classwise distribution of the following subsets can be found in Table. 2.
Subset Bots Humans
Training 25566 25528
Validation 5443 5506
Testing 5487 5462
Excess 0 6142
Table 2: Distribution of Accounts: The distribution of account types across the training, validation, testing and excess datasets. The datasets used at training are split such that the number of human and bot accounts are approximately equal, removing potential biases towards a singular class.
The Adam optimizer is used, along with a Cosine Annealing learning rate scheduler with an initial learning rate of $5\times 10^{-4}$. We deploy early stopping, defining the convergence when the validation loss is no longer decreasing after five epochs. The number of epochs for early stopping is chosen to reflect the stability of the training and account for fluctuations. Information regarding training is summarized in Table. 3. Note that the DNN is trained under the same conditions for fair comparison, modulo certain irrelevant components such as the computation of samples for learning the aleatoric uncertainty from data.
@@ 4 Datasets and Experimental Setup
5 Results
In this section, we discuss the results of predicting the labels (bot or human) of accounts with our network, a BNN inspired by the Event-Level-Uncertainty Quantification (ELUQuant) Fanelli and Giroux (2024) work originally developed for nuclear physics. This approach allows for the calculation of aleatoric and epistemic uncertainty in the predictions of whether Twitter/X accounts are bots or humans.
It is also important to compare the performance of the BNN with its deterministic counterpart, the Deep Neural Network (DNN). A BNN should perform at least as well as its deterministic counterpart. This has been demonstrated in the following way. We evaluate our model using standard methods, namely the Receiver Operating Characteristic (ROC) Curve and the associated Area Under the Curve (AUC). These metrics allow us to avoid making hard threshold cuts in the probability space and reflect a model’s performance across various thresholds. Thus, AUC is an ideal metric for such use cases. We then compare the results to the deterministic counterpart of our BNN, a DNN, along with the Random Forest (RF) from Nwala et al. Nwala et al. (2023). We extract both BLOC and Botometer features for the same set of users. Figure 4 shows a comparison between three methods, with the AUC indicated in the legend. The left plot contains the ROC curves for the algorithms trained on BLOC features, where the error is calculated through bootstrapping over the posterior on the probability. This approach allows us to produce $5\sigma$ bands over both the True Positive Rate (TPR) and False Positive Rate (FPR). The right plot provides the same ROC Curves for algorithms trained on Botometer features. We note an AUC for the BNN of $0.966\pm 0.001$, which agrees with the deterministic DNN that achieves an AUC of $0.969$ on BLOC features. The same agreement between the BNN and DNN is also seen in Botometer features with the BNN obtaining an AUC of $0.973\pm 0.001$ and the DNN obtaining an AUC of $0.975$. In both cases, the RF outperforms the DNN and BNN due to its ability to operate more efficiently with smaller training sample sizes.
(a) BLOC
(b) Botometer
Figure 4: Overlayed ROC Curves: Receiver Operating Characteristic (ROC) Curves for the Bayesian Neural Network (BNN), Deep Neural Network (DNN) and Random Forest (RF), trained on BLOC features (a) and Botometer features (b). The uncertainty band on the BNN curves is obtained through a bootstrapping method, in which we sample the posterior over the weights to obtain uncertainties on the False Positive Rate (FPR) and True Positive Rate (TPR) at each threshold. Note the DNN and BNN perform consistently within error. RF outperforms the networks due to its increased ability to operate datasets with lower statistics more efficiently.
(a) BLOC
(b) Botometer
Figure 5: Uncertainty as a Function of Probability: Epistemic uncertainty, aleatoric uncertainty and the two in quadrature as function of model probability for the models trained on (a) BLOC features (top row) and (b) Botometer features (bottom row). Note the parabolic like shape of the epistemic distribution, with maximum uncertainty around the decision boundary ($<p_{bot}>=0.5$). For a well calibrated Bayesian model, this is the expected behavior of the epistemic uncertainty. The aleatoric uncertainty is dictated by the available data, and therefore there exists no expectation on its distribution. The two uncertainties in quadrature produce a convolution of the two, epistemic and aleatoric.
We consolidated the uncertainty quantification by running a closure test, consisting of training the network (i) without including the aleatoric term, thereby predicting only the epistemic uncertainty, and (ii) including both aleatoric and epistemic terms. Details on how these two scenarios can be implemented are discussed in the Sec. 3 and are also described in Fanelli and Giroux (2024). We demonstrated through a Z-score test, visualized in Fig. 6, that the epistemic uncertainties obtained on the accounts (considering the results from methods (i) and (ii) at the account level) produce consistent epistemic uncertainties (the majority of values are within $|\text{Z}|\leq 0.5$). This closure test supports the fact that the quantified aleatoric uncertainty is decoupled from the epistemic uncertainty, with the latter appearing, on average, different from the former.
(a) BLOC
(b) Botometer
Figure 6: Z-Score Tests for Decoupling Aleatoric and Epistemic Uncertainties: Z-Score tests for the decoupling of aleatoric and epistemic uncertainties, for BLOC (a) and Botometer features (b). The results support the ability of the network to decouple epistemic and aleatoric uncertainty during training, in which the majority of samples lie within $|Z|<0.5$.
After consolidating our results, as discussed, we compare the performance of our network with respect to state-of-the-art works utilizing BLOC features Nwala et al. (2023) and Botometer Yang et al. (2022) with random forest (RF) as a classifier, as done in those papers. We evaluate the performance using precision, recall, and F1 metrics. Our method outperforms other methods in terms of recall and F1, and is nearly on par in terms of precision, as shown in Table 5. We use a threshold value of 0.5, representing the natural decision boundary of a Sigmoid function.
BLOC Botometer
Subset Model Precision Recall F1 Precision Recall F1
Human (Test Split) BNN 0.928 $\pm$ 0.003 0.921 $\pm$ 0.004 0.924 $\pm$ 0.004 0.938 $\pm$ 0.003 0.939 $\pm$ 0.003 0.939 $\pm$ 0.003
DNN 0.919 $\pm$ 0.004 0.906 $\pm$ 0.004 0.912 $\pm$ 0.004 0.927 $\pm$ 0.004 0.926 $\pm$ 0.004 0.926 $\pm$ 0.004
RF Nwala et al. (2022) 0.916 $\pm$ 0.004 0.937 $\pm$ 0.003 0.927 $\pm$ 0.004 0.938 $\pm$ 0.003 0.950 $\pm$ 0.003 0.944 $\pm$ 0.003
Bot (Test Split) BNN 0.922 $\pm$ 0.004 0.929 $\pm$ 0.003 0.925 $\pm$ 0.004 0.939 $\pm$ 0.003 0.938 $\pm$ 0.003 0.939 $\pm$ 0.003
DNN 0.908 $\pm$ 0.004 0.920 $\pm$ 0.004 0.914 $\pm$ 0.004 0.926 $\pm$ 0.004 0.927 $\pm$ 0.004 0.927 $\pm$ 0.004
RF Nwala et al. (2022) 0.936 $\pm$ 0.003 0.915 $\pm$ 0.004 0.925 $\pm$ 0.004 0.949 $\pm$ 0.003 0.938 $\pm$ 0.003 0.943 $\pm$ 0.003
Excess Human BNN N/A 0.920 $\pm$ 0.003 N/A N/A 0.939 $\pm$ 0.003 N/A
DNN N/A 0.438 $\pm$ 0.006 N/A N/A 0.501 $\pm$ 0.006 N/A
RF Nwala et al. (2022) N/A 0.938 $\pm$ 0.003 N/A N/A 0.952 $\pm$ 0.003 N/A
Table 5: Performance Comparison of Classifiers Using BLOC and Botometer Features: Bayesian Neural Network (BNN) used in our work, which is compared to a Deep Neural Network (DNN) and to a Random Forest (RF) with features from the Behavioral Language for Online Classification (BLOC) and Botometer. Precision, recall and F1 have been computed for human accounts and bot accounts (using the test dataset which is a mixture of human and bot accounts), and using other bot accounts only not present in our test dataset to test generalization.
[…]
$|P_{pred}-0.5|>3\sigma(P_{pred}).$ (9)
Equation (9) is used for $\sigma(P_{\text{pred}})=\sigma_{\text{epi.}}(P_{\text{pred}})$, $\sigma_{\text{alea.}}(P_{\text{pred}})$ and $\sigma_{\text{tot.}}(P_{\text{pred}})$, representing the cases of epistemic or aleatoric only, and the total uncertainty in quadrature. The results show an improvement in performance, across all metrics, over the baseline (i.e., the BNN without uncertainty information) as uncertainty is introduced into the decision making process. We also report the “rejection” fraction, which corresponds to the number of account that are held for further information to be acquired before classification. With regard to cuts utilizing only the epistemic component, we expect that under a robustly characterized epistemic uncertainty, withholding accounts with high epistemic uncertainty should induce performance increases being that we remove regions of the feature space where overlap (between human and bot accounts) persists to a high degree, i.e., regions of low confidence or higher uncertainty. Similarly, with regard to the cuts using only the aleatoric component, we expect that withholding accounts with high uncertainty should induce performance increases being that we remove regions of high stochasticity in the feature space,corresponding to potentially unreliable predictions due to lack of information seen at training time, or simply regions that are well defined but have high variance. The usage of these two in quadrature can then further increase performance by addressing both issues in unison at inference. This additional information from uncertainty allows for more informed decisions, thereby impacting the decision-making process for Twitter/X accounts, as further discussed below.
Human Bot
BLOC Baseline 92.5 $\pm$ 0.3 0.928 $\pm$ 0.004 0.921 $\pm$ 0.004 0.924 $\pm$ 0.004 0 0.922 $\pm$ 0.004 0.929 $\pm$ 0.004 0.925 $\pm$ 0.004 0
Epistemic 95.8 $\pm$ 0.2 0.958 $\pm$ 0.003 0.958 $\pm$ 0.003 0.958 $\pm$ 0.003 10.6 0.959 $\pm$ 0.003 0.959 $\pm$ 0.003 0.959 $\pm$ 0.003 12.2
Aleatoric 96.0 $\pm$ 0.2 0.959 $\pm$ 0.003 0.961 $\pm$ 0.003 0.960 $\pm$ 0.003 12.2 0.962 $\pm$ 0.003 0.960 $\pm$ 0.003 0.961 $\pm$ 0.003 14.0
Quadrature 96.6 $\pm$ 0.2 0.964 $\pm$ 0.003 0.967 $\pm$ 0.003 0.966 $\pm$ 0.003 15.5 0.968 $\pm$ 0.003 0.965 $\pm$ 0.003 0.967 $\pm$ 0.003 17.6
Botometer Baseline 93.9 $\pm$ 0.2 0.938 $\pm$ 0.003 0.939 $\pm$ 0.003 0.965 $\pm$ 0.003 0 0.939 $\pm$ 0.003 0.938 $\pm$ 0.003 0.939 $\pm$ 0.003 0
Epistemic 96.5 $\pm$ 0.2 0.962 $\pm$ 0.003 0.968 $\pm$ 0.003 0.965 $\pm$ 0.003 7.3 0.968 $\pm$ 0.003 0.962 $\pm$ 0.003 0.965 $\pm$ 0.003 7.5
Aleatoric 96.5 $\pm$ 0.2 0.962 $\pm$ 0.003 0.968 $\pm$ 0.003 0.965 $\pm$ 0.003 7.8 0.969 $\pm$ 0.003 0.963 $\pm$ 0.003 0.966 $\pm$ 0.003 10.0
Quadrature 97.2 $\pm$ 0.2 0.968 $\pm$ 0.003 0.975 $\pm$ 0.003 0.972 $\pm$ 0.003 10.3 0.976 $\pm$ 0.003 0.968 $\pm$ 0.003 0.972 $\pm$ 0.003 12.1
Table 6: Performance Comparison with $3\sigma$ Uncertainty Thresholds: Evaluation of the uncertainty informed decision making process, in which a $3\sigma$ cut is applied on the probability around 0.5 to indicate whether decisions should be made about a user account. $3\sigma$ cuts are applied to the epistemic, aleatoric and uncertainties in quadrature. Note the increase in performance as specific accounts are withheld. Extra information from these accounts can be acquired to reduce uncertainty to a desirable threshold, or allow human intervention on more reasonable sample sizes.
In summary, our approach features a fully Bayesian framework inspired by ELUQuant Fanelli and Giroux (2024) to classify Twitter/X accounts as bots or humans, assessing its performance against DNNs and RF models. The BNN demonstrates comparable performance to both the DNN and RF models in terms of AUC, while providing additional information in the form of uncertainty, which is crucial for decision-making at the account level. Closure tests affirm the robustness of our uncertainty quantification, in which we are able to decouple the aleatoric and epistemic components. By applying a $3\sigma$ uncertainty threshold, we observe improved accuracy and F1 scores, highlighting the utility of uncertainty-aware models in bot detection. This approach not only enhances predictive reliability but also provides deeper insights into model behavior.
6 Conclusions